Write parametric equations of the line 2x+3y=11
Answer by Theo(13342)

solve for y to get y = (11 - 2x) / 3

let x = t and this equation becomes:

y = (11 - 2t) / 3

your parametric equation is:

x = t
y = (11 - 2t) / 3

if you plot this equation, you should get the same graph as the equation of

2x + 3y = 11